Oriole Industries carries no inventories. Its product is manufactured only when a customer’s order is received. It is then shipped immediately after it is made. For its fiscal year ended October 31, 2020, Oriole’s break-even point was $1.33 million. On sales of $1.20 million, its income statement showed a gross profit of $228,000, direct materials cost of $401,000, and direct labor costs of $505,000. The contribution margin was $192,000, and variable manufacturing overhead was $50,000.

Calculate the following:

1. Variable selling and administrative expenses.

$

2. Fixed manufacturing overhead.

$

3. Fixed selling and administrative expenses.

$
Answer #1

Break even point = $1,330,000

Sales = $1,200,000

Gross profit = $228,000

Direct materials = $401,000

Direct labor = $505,000

Contribution margin = $192,000

Variable manufacturing overhead = $50,000

Total variable expense = Sales - Contribution margin

= 1,200,000-192,000

= $1,008,000

Total variable expenses = Direct materials + Direct labor + Variable manufacturing overhead + Variable selling and administrative expense

1,008,000 = 401,000+505,000+50,000+Variable selling and administrative expense

Variable selling and administrative expense = $52,000

Contribution margin ratio = Contribution margin / Sales

= 192,000/1,200,000

= 16%

Break even point = Total fixed cost / Contribution margin ratio

1,330,000 = Total fixed cost/ 16%

Total fixed cost = $212,800

Cost of goods sold = Sales - Gross profit

= 1,200,000-228,000

= $972,000

Cost of goods sold = direct materials + Direct labor + Variable manufacturing overhead + Fixed manufacturing overhead

972,000 = 401,000+505,000+50,000+Fixed manufacturing overhead

Fixed manufacturing overhead = $16,000

Total fixed expense = Fixed manufacturing overhead + Fixed selling and administrative expense

212,800 = 16,000 + Fixed selling and administrative expense

Fixed selling and administrative expense = $196,800

1. Variable selling and administrative expenses.

$52,000

2. Fixed manufacturing overhead.

$16,000

3. Fixed selling and administrative expenses.

$196,800

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks